# Reading Subtext: Evaluating Large Language Models on Short Story Summarization with Writers

### v1

@@ 6 Results
We first discuss the findings from the quantitative metrics, and then review our error analysis and the qualitative feedback from the writers.
@@ 6.1 How good are the summaries?
Interestingly, we find that while models have the appearance of fluency, they are given a less than perfect score for coherence 40% or more of the time as there is more to a coherent narrative summary than just fluency. In the opposite direction, the models do surprisingly well at analysis. The writers felt that 56% of the GPT-4 summaries summarized some of the themes and interpretation they had hoped to communicate as a writer, which is a challenging task even for humans (see examples of the best analysis in Figure 4).
@@
On the worst summaries, they commented things like: "completely misses the fundamental thread", "tries to make too many analyses like a high school english class while missing the overall bigger feelings", and "makes use of…stock phrases or stand-ins rather than accurate, specific summary".
Also, for one of the stories Claude refused to summarize, the writer commented, "Art is such an important tool for processing. I don’t like that the summary wasn’t able to process a recurring vague nightmare."
@@ 6.2 What mistakes do the models make?
Narrative Element
GPT-4 Claude Llama Avg.
Feeling
45% 20% 15% 27%
Causation
20% 20% 10% 17%
Action
5% 20% 25% 17%
Character
5% 5% 30% 13%
Setting
15% 15% 15% 15%
No-consensus 10% 20% 5% 12%
Table 4: Percent of faithfulness errors categorized under each narrative element for each model.
Figure 6: Examples of faithfulness errors in each of the categories assigned by the annotators.
In Figures 6 and 7, we show examples of issues models have in each attribute – coverage, faithfulness, coherence, and analysis. In general, we find that while writers rank GPT-4 and Claude summaries somewhat equally, Claude summaries often leave out important details that lower their coverage score and create issues with faithfulness as identified by the numerical scores. For example, while the GPT-4 coverage error in Figure 7 oversimplifies a major plot point in the story, the Claude error removes an entire thematic subplot and creates a faithfulness error that the mother is still ill instead of recovered.
By categorizing the faithfulness errors in Table 4, we find that most faithfulness errors are in Feeling when averaged across the models. These errors require interpretation of character reactions and internal states, so this result demonstrates models still struggle with reading people. However, the error distribution for each model varies. For example, Llama has a high percentage of Character errors. This is likely an artifact of the chunk-then-summarize strategy as sometimes characters are conflated or split in two as the model fails to track their details across different chunk summaries (see Character example in Figure 6). Chang et al. (2023) also found a similar type of error (entity omission) to be most common with chunk-then-summarize approaches.
@@
Lastly, some of the writers commented that the models would copy direct and unique wording from the stories without placing it in quotations, which bordered on plagiarism. We flag this for future work.
Figure 7: Examples of coverage, coherence, and analysis errors the models make.
@@ 6.3.1 Length
Figure 8: Averaged writer scores for summaries for stories within a given length bucket for each model. The bar saturation indicates the length bucket.
In Figure 8, we compare the average scores across different story length buckets. The stories in the short bucket are short enough that Llama can fit the entire story in its input context. Looking at this bucket, the scores are better for Llama than the longer buckets but still not better than the other two models. For Claude and GPT-4, it seems the long-context models can summarize short and long stories equally well (up to 10,000 tokens). For these two models, the summaries of the longest stories have the highest average scores by a slim margin, but this may be because there are fewer stories in this bucket (only 6).
@@ 6.3.2 Reading-Level
We plot the reading-level (estimated years of education needed to understand the wording easily) against the writer-assigned scores averaged across the four attributes for each summary. We find no significant correlation between story reading-level and writer-assigned summary scores for any of the models (GPT-4: $r$11, $p$61; Claude: $r$33, $p$12, Llama: $r$23, $p$27). The average Flesch-Kincaid grade-level for the group of stories is 6th grade. As you can see in the 2nd grade example in Figure 3 though, which features someone being shot, this score captures an aspect of the wording but does not consider age-appropriate content or conceptual complexity.
@@ 6.3.3 Narrating, Story, and Discourse
In Table 5, we see that stories with unreliable narrators are harder for all three models to summarize. All three models have lower average scores in this bucket. As an example of an error due to an unreliable narrator, the same unreliable narrator who is quoted in Section 5.2 further describes himself as wanting a "small, normal life", and then describes his new love interest:"I met a girl… I think she’s just like me. She loves traveling… It’s pretty cool." GPT-4 mistakenly characterizes this interaction as: "The protagonist meets [a girl]… with whom he shares… a desire for a simple life." even though there is no evidence she likes a simple life (or even that we should believe the narrator that he does), other than that the protagonist has described her as "like [him]".
@@
For story and discourse, the models have mixed performance across the different buckets. Llama does have lower scores in the buckets we hypothesized might be more challenging. Overall, it seems that unreliable narrators are a challenge for LLMs, whereas level of detail and flashbacks depend on the case. One writer commented directly on the issue of an unreliable narrator by observing, "there’s a second layer here that the summary misses somewhat—the narrator is not really to be trusted".
6.3.4 Can an LLM replace the writers in evaluating the summaries?
@@
Figure 9: Confusion between model-assigned scores/labels and human-assigned scores/labels for GPT-4 and Claude.
Attribute GPT-4 Claude
Summary Rating: Coverage .46* .18
Faithfulness .37* .05
Coherence .18 .02
Analysis .21* .02
Faithfulness Error Labels .65 .53
Table 6: Agreement between model-assigned scores/labels and writer-assigned scores/labels. For the summary ratings which are Likert scores, we report Pearson’s $r$ (* indicates p-value <.05). For the faithfulness errors which are categorical, we report Cohen’s Kappa.
@@
In Figure 9, we compare the GPT-4 and Claude scores to the writer scores from Table 3, and we see considerable confusion across the ratings (first row of plots). Claude mostly only rates summaries as 3 or 4, causing overestimation of ratings (3.67 average score vs. 3.01 from the writers).
@@
In terms of attributes, both models rate the coherence of the summaries much higher than the writers do (average scores: GPT-4 - 3.84, Claude - 3.82, Writers - 3.33). GPT-4 largely underestimates the coverage score for all three models (average score 2.68 vs. 3.01 from the writers). Many of the writers commented that the GPT-4 summaries were too detailed, so the model may struggle with identifying what is most important to cover. Claude overestimates performance on faithfulness and analysis. It gives 85% of summaries a 4 on faithfulness on average and 85% on analysis as well, relative to averages 27% and 39% respectively from the writers. GPT-4 also overestimates faithfulness performance just for itself (64% scores of 4 vs. 44% from the writers).
@@
We also show the Pearson’s $r$ correlation between the model scores and the writers’ scores in Table 6 across attributes. We see no correlation for Claude scores and some statistically significant but weak correlation for GPT-4 scores, other than for coverage which has moderate positive correlation. We also find no significant correlation between BooookScore coherence scores and the writer coherence scores ($r$:.11, $p$:.18). The average BooookScore scores are almost the same across all three models (GPT-4: .95, Claude: .93, Llama: .93), which does not capture the significant difference in coherence that the writers judged in Llama in particular. Overall, these results support prior work Chakrabarty et al. (2023a), which shows that LLMs are not yet reliable evaluators for skilled writing tasks.
However, we do find that GPT-4 and Claude have moderate to substantial agreement with our human labelers for categorizing the faithfulness errors (see Table 6 and the second row in Figure 9). Importantly, they are not actually detecting the faithfulness errors themselves, but performing this error categorization could be helpful for future error analysis in studies like this.
@@ 7 Discussion
Our results show that models are capable of thorough and interesting analysis of the themes present in unseen stories. When models are asked to summarize stories that may have analysis available online, it isn’t clear if the models are simply regurgitating analysis from the training data. Therefore, it is important to examine their capabilities on original and challenging content. Additionally, while we expected it might be challenging for LLMs to identify salient information within long stories, we find that long-context models demonstrate as good understanding of longer stories as shorter ones.
The models also have notable weaknesses. We demonstrate that even the best models are making significant errors across all summary attributes on about half of summaries. Many of the faithfulness errors come from misinterpreting challenging subtext and particularly character feelings and reactions. This will be an important area for future study as it demonstrates language models still struggle with aspects of theory of mind [^10] Theory of mind is the ”ability to understand and keep track of the mental states of others” Xu et al. (2024)., one of the most challenging being understanding an unreliable narrator. Additionally, our analysis of LLM-based evaluation shows that while LLMs can assist with evaluation, they should not replace the expert human judgments.
@@ 8 Conclusion
We work with writers to provide unpublished short stories and evaluate the quality of LLM-generated summaries of these stories. We present a quantitative and qualitative holistic evaluation of summary quality grounded in narrative theory that is based on data LLMs did not train on. In doing this, we identify that LLMs can demonstrate understanding of long narrative and thematic analysis, but the remaining challenges lie in reliable interpretation of subtext - specifically feeling and narrative voice. Our methodology sets an important example of how we can collaborate with domain experts to reach beyond the paradigm of evaluating LLMs with LLMs on data they may have been trained on.
@@ Appendix A Information Sheet for Informed Consent
Voluntary Participation: You may stop participating at any time without penalty by emailing {PI} ({PI-email}). At that time, we will delete any short stories you may have shared with us. You will only be compensated for stories for which you complete the full evaluation process. We may end your participation if you are not following the instructions.
@@
Questions/concerns: You may email questions to the principal investigator, {PI}: {PI-email}. If you feel you have been treated unfairly you may contact the Columbia University Institutional Review Board and reference Protocol AAAU8875. Please respond to this email by copying the below consent paragraph to indicate that you understand the information in this consent form. You have not waived any legal rights you otherwise would have as a participant in a research study. If you have any questions about your rights or responsibilities as a research participant, please contact the Columbia University HRPO office at: Phone {phone-number}; Email {email}. We recommend either printing this information sheet or taking note of this contact information before beginning the study in case of questions.
@@ B.1 GPT-4
Faithfulness Error Labeling:
System prompt
{instructions loaded from same file as for human annotators}
[…]
@@ B.2 Claude
Prompt (questions asked one at a time in this format)
[…]
{AI_PROMPT}
Faithfulness Error Labeling:
@@ Appendix C Short Story Uploader and Evaluation Interfaces
Appendix D Faithfulness Error Annotation Instructions
@@
Figure 11: Annotator instructions and an example error for annotation of faithfulness error labels.

### v3

@@ 6.1 How good are the summaries?
Figure 4: Distribution of Likert score ratings for each model’s summaries by attribute.
@@
Interestingly, we find that while models have the appearance of fluency, they are given a less than perfect score for coherence 40% or more of the time as there is more to a coherent narrative summary than just fluency. For analysis, writers felt that 56% of the GPT-4 summaries summarized some of the themes and interpretation they had hoped to communicate as a writer, which is a challenging task even for humans (see examples of the best analysis in Figure 5). However, in Table 3, we see that summaries rated 4 for analysis actually contain at least one error in analysis on average. Furthermore, by breakdown of span-level error counts in Table 5, we see that writers find as many errors in analysis as in faithfulness. These numbers suggest that while writers may have been impressed by the models’ ability to do any analysis in their ratings, analysis remains a challenge.
@@
On the worst summaries, they commented things like: "completely misses the fundamental thread", "tries to make too many analyses like a high school english class while missing the overall bigger feelings", and "makes use of…stock phrases or stand-ins rather than accurate, specific summary". Also, for one of the stories Claude refused to summarize, the writer commented, "Art is such an important tool for processing. I don’t like that the summary wasn’t able to process a recurring vague nightmare."
@@ 6.2 What mistakes do the models make?
Error Type GPT4 Claude Llama Tot.
Coverage:
insignificant 4 2 8 14
vague 15 19 30 64
Sub-Total 19 21 38 78
Analysis:
unsupported 18 16 70 104
Coherence:
repetition 3 4 0 7
missing context 11 11 15 37
inconsistent 1 3 1 5
abrupt transition 3 6 8 17
Sub-Total 18 24 24 66
Faithfulness:
feeling 5 8 10 23
Causation 2 9 4 15
Action 5 6 28 39
Character 2 1 15 18
Setting 2 2 5 9
Sub-Total 16 26 62 104
Total 71 87 194
Table 5: Number of errors categorized under each narrative element by the writers for each model.
Figure 7: Examples of span-level errors assigned by the writers in each of the categories.
In Figure 7, we show examples of issues models have in each attribute. In general, we find that while writers rank GPT-4 and Claude summaries somewhat equally, Claude summaries often leave out important details that lower their coverage score and create issues with faithfulness as identified by the numerical scores. For example, while the GPT-4 coverage error in Figure 7 includes too many details, for one story, Claude leaves out an entire thematically important subplot. The narrator desperately needs money, and Claude fails to mention she is in debt due to a medical crisis.
Looking at the categorization of errors in Table 5, we find that vagueness is a significant issue across all three models. We see there is a large amount of unsupported analysis for all three models, and this is one of the biggest error categories overall. In coherence, the biggest error category is missing context. This overlaps with coverage and lets us know that significant details are left out or too vague to fully understand. Lastly, we find that most faithfulness errors across the models are in Action and Feeling. These errors require interpretation of what characters did and how they reacted or felt.
For example, in Figure 7, the Feeling error is a misinterpretation of who the narrator is attracted to. Claude fails to interpret phrases like "He looked nice", and "I felt warm in the cheeks", as signs of attraction to Arkady. Another feeling error describes a character as having a difficult relationship with their mother, when actually they are quite close, but the circumstances around their relationship have been difficult because the mother almost died from cancer. One writer comments on another summary, "The ending of contentment is supposed to be contemplation. [The summary] leaves out any feeling of hesitancy or reflection". The action error example in Figure 7 mistakenly states that the narrator decided to stay at the summit. However, "The cold settled into me", is meant to imply that the narrator died at the summit, which was not an action or choice.
Taken in conjuction, all of these results show that models struggle with identifying the right level of specificity and with interpreting subtext. Three writers commented on subtext specifically, saying, "Subtext is missing entirely" on a Llama summary, "Everything that’s ‘text’ is accurately represented and well summarized, but the ‘subtext’ is either incorrectly interpreted or missing" on a Claude summary, and "The summary struggles with subtext" on a GPT-4 summary.
However, the error distribution for each model varies. For example, Llama has a high percentage of Character errors. This is likely an artifact of the chunk-then-summarize strategy as sometimes characters are conflated or split in two as the model fails to track their details across different chunk summaries (see Character example in Figure 7). Chang et al. (2024) also found a similar type of error (entity omission) to be most common with chunk-then-summarize approaches.
@@
Lastly, some of the writers commented that the models would copy unique wording from the stories without placing it in quotations, which bordered on plagiarism. We flag this for future work and include statistics on it in Table 2.
@@ 6.3.1 Length
Figure 8: We plot story length (in tokens) and summary length (in word count) against writer summary ratings averaged across the four attributes. In the left plot, we show the line of best fit and Pearson’s $r$ with $p$-value for each model individually. In the right plot, we show the correlation across the full set of summaries.
In Figure 8, we look at the correlation between story length and summary rating (left plot). We do see a downward trend for Llama. Llama’s summaries get worse as stories get longer, which is expected given the chunk-then-summarize method it employs for longer stories. However, the correlation is not quite statistically significant. For Claude and GPT-4, it seems the long-context models can summarize short and long stories equally well (up to 10,000 tokens).
In Figure 8, we also compare the average scores across different summary lengths (right plot), to check if there is a bias in ratings to shorter or longer summaries. We do not find any correlation between summary length and quality.
@@ 6.3.2 Reading-Level
We plot the reading-level (estimated years of education needed to understand the wording easily) against the writer-assigned scores averaged across the four attributes for each summary. We find no significant correlation between story reading-level and writer-assigned summary scores for any of the models (GPT-4: $r$11, $p$61; Claude: $r$33, $p$12, Llama: $r$23, $p$27). The average Flesch-Kincaid grade-level for the group of stories is 6th grade. As can be seen in the 2nd grade example in Figure 3 though, which features someone being shot, this score captures an aspect of the wording but does not consider age-appropriate content or conceptual complexity.
@@ 6.3.3 Narrating, Story, and Discourse
In Table 6, we see that stories with unreliable narrators are harder for all three models to summarize. All three models have lower average scores in this bucket. As an example of an error due to an unreliable narrator, the same unreliable narrator who is quoted in Section 5.3, further describes himself as wanting a "small, normal life", and then describes his new love interest: "I met a girl… I think she’s just like me. She loves traveling… It’s pretty cool." GPT-4 mistakenly characterizes this interaction as: "The protagonist meets [a girl]… with whom he shares… a desire for a simple life." even though there is no evidence she likes a simple life (or even that we should believe the narrator that he does), other than that the protagonist has described her as "like [him]".
@@
For story and discourse, the models have mixed performance across the different buckets. Llama does have lower scores in the buckets we hypothesized might be more challenging. Overall, it seems that unreliable narrators are a challenge for LLMs, whereas level of detail and flashbacks depend on the model. One writer commented directly on the issue of an unreliable narrator by observing, "there’s a second layer here that the summary misses somewhat—the narrator is not really to be trusted".
6.4 Can an LLM replace the writers in evaluating the summaries?
@@
Figure 9: Confusion between model-assigned scores and human-assigned scores for GPT-4 and Claude.
LLM ROUGE BERTScore Lla.2 Faith. Metrics
GPT4 Clau. BkS Fab. R1 R2 RL RLS Prec. Rec. F1 PPL AlS UnE MiC
Cov. .$46^{*}.$18 - - .$33^{*}.$$36^{*}.$$32^{*}.$$30^{*}.$$29^{*}.$$21$$25^{*}.$15 - - -
Fai. .$37^{*}.$05 - .$29^{*}.$$18$$28^{*}.$$18$$17$$27^{*}.$$25^{*}.$$27^{*}.$00 -.04 .$32^{*}.$$20^{*}.$
Coh. .18 .02 .11 - .$18$$21$$16$$16$$29^{*}.$.$01$$09 -.$
Ana. .$21^{*}.$02 - - -.$05$$01$$06$$08$$15$$14$$16$12 - - -
Avg. - - - - .$21$$28^{*}.$$20$$18$$32^{*}.$$20$$26^{*}.$14 - - -
Table 7: Correlation between automatic metrics and writer-assigned scores. We report Pearson’s $r$ (* indicates p-value <.05). The ‘Avg.’ row takes the average score across the four attributes as the writer score. We do not report correlation for attributes the metrics are not designed for. Metric Key for uncommon abbreviations: BkS - BooookScore, Fab. - FABLES, RLS - ROUGE-LSum, AlS - AlignScore, UnE - UniEval, MiC - MiniCheck.
@@
In Figure 9, we compare the GPT-4 and Claude scores to the writer scores from Table 4, and we see considerable confusion across the ratings. Claude mostly only rates summaries as 3 or 4, causing overestimation of ratings (3.67 average score vs. 3.01 from the writers).
@@
In terms of attributes, both models rate the coherence of the summaries much higher than the writers do (average scores: GPT-4 - 3.84, Claude - 3.82, Writers - 3.33). GPT-4 largely underestimates the coverage score for all three models (average score 2.68 vs. 3.01 from the writers). Many of the writers commented that the GPT-4 summaries were too detailed, so the model may struggle with identifying what is most important to cover. Claude overestimates performance on faithfulness and analysis. It gives 85% of summaries a 4 on faithfulness on average and 85% on analysis, relative to averages of 27% and 39% respectively from the writers. GPT-4 also overestimates faithfulness performance just for itself (64% scores of 4 vs. 44% from the writers).
@@
We show the Pearson’s $r$ correlation between the model scores and the writers’ scores in Table 7 across attributes. We see no correlation for Claude scores and some statistically significant but weak correlation for GPT-4 scores (indicating there may be a relationship but the scores are not well calibrated). Coverage is the one attribute with moderate positive correlation. We find no significant correlation between BooookScore coherence scores and the writer coherence scores ($r$:.11, $p$:.18). The average BooookScore scores are almost the same across all three models (GPT-4: .95, Claude: .93, Llama: .93), which does not capture the significant difference in coherence that the writers judged in Llama in particular. We find significant correlation between FABLES and the writer faithfulness scores but the correlation is weak and less than just using GPT-4 ($r$: .29, $p$: .01). The average FABLES scores are also approximately the same across the three models (GPT-4: .99, Claude: .99, Llama: .99). Overall, these results support prior work Chakrabarty et al. (2024), which shows that LLMs are not yet reliable evaluators for skilled writing tasks.
For context, we also include simple automatic metrics and recent Faithfulness-specific finetuned models. The only metric which performs better than GPT-4 is BERTScore for coherence using the story as a reference. However, the correlation is weak, and Goyal et al. (2022c) has already shown that BERTScore does not penalize many coherence errors. We additionally include the Llama-2-7B computed perplexity scores in this correlation table to check if there is any relationship between the similarity of stories to LLM training data and attribute ratings, and we find none.
@@ 7 Discussion
Our results show that, in the best summaries, models are capable of some interesting analysis of the themes present in unseen stories. When models are asked to summarize stories that may have analysis available online, it is not clear if the models are simply regurgitating analysis from the training data. Therefore, it is important to examine their capabilities on original and challenging content. Additionally, while we expected it might be challenging for LLMs to identify salient information within long stories, we find that long-context models demonstrate as good understanding of longer stories as shorter ones.
The models also have notable weaknesses. We demonstrate that even the best models are making significant errors across all summary attributes on about half of summaries. They are often too vague or missing important context, and they struggle with challenging subtext, providing unsupported analysis and misinterpreting character feelings/reactions and actions. This will be an important area for future study as it demonstrates language models still struggle with aspects of theory of mind [^10] Theory of mind is the ”ability to understand and keep track of the mental states of others” Xu et al. (2024)., one of the most challenging being understanding an unreliable narrator. Additionally, our analysis of LLM-based evaluation shows that they should not replace the expert human judgments.
@@ 8 Conclusion
We work with writers to provide unpublished short stories and evaluate the quality of LLM-generated summaries of these stories. We present a holistic evaluation at the span, summary, and story level of summary quality grounded in narrative theory that is based on data LLMs did not train on. We identify that LLMs can demonstrate understanding of long narrative and thematic analysis, but they struggle with specificity and reliable interpretation of subtext and narrative voice. Our methodology sets an important example of how we can collaborate with domain experts to reach beyond the paradigm of evaluating LLMs with LLMs on data they may have been trained on.
Acknowledgments
We would like to express our gratitude to the short story writers for sharing their work and contributing annotations. Additionally, we would like to thank our reviewers and Action Editor for their thoughtful feedback.
@@ Appendix A Information Sheet for Informed Consent
Voluntary Participation: You may stop participating at any time without penalty by emailing Kathleen McKeown (kathy@cs.columbia.edu). At that time, we will delete any short stories you may have shared with us. You will only be compensated for stories for which you complete the full evaluation process. We may end your participation if you are not following the instructions.
@@
Questions/concerns: You may email questions to the principal investigator, Kathleen McKeown: kathy@cs.columbia.edu. If you feel you have been treated unfairly you may contact the Columbia University Institutional Review Board and reference Protocol AAAU8875. Please respond to this email by copying the below consent paragraph to indicate that you understand the information in this consent form. You have not waived any legal rights you otherwise would have as a participant in a research study. If you have any questions about your rights or responsibilities as a research participant, please contact the Columbia University HRPO office at: Phone {phone-number}; Email {email}. We recommend either printing this information sheet or taking note of this contact information before beginning the study in case of questions.
@@ Appendix C Short Story Uploader and Evaluation Interfaces
Appendix D Fine-Grained Error Annotation Interface
@@
Cost: \$1,150 to compensate writers for their work
Figure 11: Top panel shows the instructions for the fine-grained error annotation task. Bottom panel shows the interface which was created using Doccano Daudert (2020). A user can highlight a span of text and then select a label from the dropdown to assign to it.